OpenStudy (anonymous):

a horse canters away from a trainer in a straight line, moving 116m away in 14s. It turns abruptly and gallops halfway back in 4.8s. What's its average velocity for the entire trip and using 'away from trainer' as positive direction?

OpenStudy (anonymous):

At the end of its chaotic gallop, the horse is 58m away from its starting point. Therefore its displacement=58m And it took a total of 18.8s to come to its senses. Therefore, time taken=18.8s So, average velocity=displacement/time taken =58m/18.8s =3.08 m/s.

OpenStudy (anonymous):

thank you, i had calculated the displacement wrong.. i took x1= 116 and x2=-58 :/ can i ask another one please?

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

thank you!^^ a car traveling 95km/h is 110m behind a truck traveling 75km/h, how long will it take the car to reach the tuck?

OpenStudy (anonymous):

first of all, the relative velocity of the car w.r.t the truck=95-75 =20km/hr =20*1000/3600 =5.55m/s

OpenStudy (anonymous):

so, the time it takes to travel 110m is t=distance/speed =110/5.5 =20s

OpenStudy (anonymous):

thank you!